what is the distance between any two consecutive layers in ccp arrangement and how to calculate it?

Grade:12

1 Answers

sree lakshmi
4 years ago
in a ccp fashion  the vertical distance between any two consecutive layer is = a
in ccp a2= 4r
 a= 4r/2
 a3  =4r \left ( \sqrt{3} \right )/{\sqrt{2}}